OpenStudy (anonymous):

Help me check these derivatives? A) K=e^3r+r^5e^-.09r My answer: 3e^3r+5r^4(e^-0.9r)*(-0.9e^-0.9r) B) G(x)=15(0.57)^x/6 My answer: (8.55)^x/6*(.57x/6)^1-x/6

OpenStudy (turingtest):

I think you forgot a factor of r^5 in the product rule in your first answer

OpenStudy (turingtest):

for the second one I think you used the wrong rule

OpenStudy (anonymous):

I used chain rule for the 2nd and first. Could you show me how you would solve them?

OpenStudy (turingtest):

I'm sorry I have to go, but someone else will help you.

OpenStudy (gorv):

how do you apply chain rule ??

OpenStudy (gorv):

\[r^5*e^{-0.9 r}\]

OpenStudy (anonymous):

Would I not apply chain rule for that?

OpenStudy (anonymous):

Or would I just take the derivative of each?

OpenStudy (gorv):

show me your steps

OpenStudy (anonymous):

Alright. 5r^4*(e^-.9r)*(-.9e^.-9r)

OpenStudy (anonymous):

That's what i did...

OpenStudy (anonymous):

r^5(z)*(z')

OpenStudy (gorv):

\[=U*\frac{ dV }{ dr }+V\frac{ dU }{ dr }\]

OpenStudy (anonymous):

and take the derivative of r^5, which is 5r^4

OpenStudy (gorv):

first we will keep second term as it is and derivative of first than add that with derivative of second and keep first as it is

OpenStudy (anonymous):

Why? What rule is this?

OpenStudy (gorv):

5r^4*e^-0.9r + ???

OpenStudy (anonymous):

-.9e^-.9r?

OpenStudy (gorv):

product rule

OpenStudy (anonymous):

So you're essentially saying the product rule is almost exactly like the chain rule except you add and not multiply?

OpenStudy (gorv):

you know formula for it ??

OpenStudy (anonymous):

For the product rule? I don't think so.

OpenStudy (gorv):

well in product rule we only differentiate one and keep others untouched

OpenStudy (gorv):

r^5*e^-0.9r they are two terms r^5 and e^-0.9r

OpenStudy (gorv):

choose one term which you want to differentiate ???

OpenStudy (gorv):

@hhipkiss pls reply

OpenStudy (anonymous):

So, the adding would be r^5*-.9e^-.9r?

OpenStudy (anonymous):

3e^3r+(5r^4)*(e^-.9r)+(r^5*-.9e^-.9r)?

OpenStudy (gorv):

a little mistake

OpenStudy (anonymous):

Which is?

OpenStudy (gorv):

minus sign will come out as both term are in multiplication

OpenStudy (gorv):

on second part (-r^5*9*e^-0.9r)

OpenStudy (anonymous):

oh, yeah, gotcha.

OpenStudy (anonymous):

Could you assist me with B possibly?

OpenStudy (gorv):

base is constant ..agree??

OpenStudy (anonymous):

the 15? Yes.

OpenStudy (anonymous):

I was thinking of maybe using A^x=(lna)(a^x)?

OpenStudy (gorv):

yes :)

OpenStudy (gorv):

15 is constant so keep it aside we will bring it later

OpenStudy (gorv):

(0.57)^x/6

OpenStudy (gorv):

what is a here ????

OpenStudy (anonymous):

A is .57.